Probability and Statistics II GSIA, Carnegie Mellon University 45-734, Spring 2002

Homework #3, due Tuesday, April 9, 2002

Please attach relevant eview output to your assignment!

Please do the following problems from the book:

pg 524 47

Also, please use the auto data to construct a model of auto pricing. Suppose you think price is a function of size of car (proxied by weight), fuel efficiency, and quality (proxied by consumer reports reliability)

Throughout, be clear what you are assuming.

1. Test at the 5% level the hypothesis that size does not affect price

2. Test at the 10% level the hypothesis that neither efficiency nor quality affect price

3. Make a 90% confidence interval for the effect of one extra lb on price.

4. Test at the 1% level the hypothesis that none of these factors affect price

5. Test at the 1% level that the a one point increase in reliability is equivalent to a 400lb increase in size, in terms of its effect on price.